The Johnson County Carnival costs $15 plus $0.50 per ride. The Jefferson County Carnival costs $10 plus $1.00 per ride. How many rides should you plan on riding to make the Johnson County Carnival less expensive than the Jefferson County Carnival?
Answer by mananth(16949)

Let r be the number of rides
The total cost for r rides =
Johnson- C(j)= 15+0.5r
Jefferson= C(Je)=10+1r
Johnson < Jefferson
C(j) 15+0.5r<10+1r
add -0.5r
15+0.5r-0.5r<10+1r-0.5r
15<10+0.5r
-10
15-10<0.5r
5<0.5r
/0.5
10 so r > 10